# Consistent Validation for Predictive Methods in Spatial Settings

### v2

@@ D.3.5 Inconsistency of Nearest Neighbors with Number of Neighbors Depending On Number of Validation Points
The idea is that either 1.) $(k_{n})_{n=1}^{\infty}$ has a bounded sub-sequence, in which case along this sub-sequence, the $\hat{R}_{N^{\textup{{val}}}}(h)$ can have a variance bounded below by 00, and so by Proposition D.8 these estimators are bounded away from the test risk with fixed probability. Or 2.) the number of neighbors used tends to infinity, in which case we can find a sequence of data that accumulates more slowly around the test site, leading to many neighbors far from the point being used in the estimator, and therefore non-negligible bias.
@@
$\mathrm{Var}(\hat{R}_{N^{\textup{{val}}}})$ $=\mathrm{Var}(\sum_{n=1}^{N^{\textup{{val}}}}w^{\textup{{NN}},k}|\epsilon_{n}|)$ (46)
@@
Applying Proposition D.8 to the random variable $|\hat{R}_{N^{\textup{{val}}}}(h)-R(h)|$ we conclude there exists an $\epsilon,\delta$ such that with probability at least $1-\delta$,
@@
D.4 General Nearest Neighbor Bound and Selecting the Number of Neighbors
@@
Proposition D.10 (Minimization over Powers of Two).
@@
D.4.1 Preliminary Result: Hoeffding’s Inequality
@@
Lemma D.11 (Hoeffding’s Inequality, [Hoeffding, 1963, Theorem 2]).
@@
D.4.2 Proof of General Nearest Neighbor Risk Estimation Bound
We again recall and prove Theorem 5.1. The idea is to apply triangle inequality to split the error into a bias term and a sampling error term. The sampling error term is then controlled with Lemma D.11 since the loss is bounded. The bias term is controlled using Assumption 2.5. See 5.1
@@ D.4.2 Proof of General Nearest Neighbor Risk Estimation Bound
The first term, $\tau_{1}$ is a bias term, while the second term, $\tau_{2}$ is a sum of $N^{\textup{{val}}}$ independent variables with expectation zero. Using Assumption 2.1, we apply Hoeffding’s inequality (Lemma D.11) to bound $\tau_{2}$: for any $\delta\in(0,1)$ with probability at least $1-\delta$
@@
Substituting this in Equation 61 and using Equation 60
@@
The result follows from combining Equations 58, 59, 70 and 64. ∎
D.4.3 Proofs Related to Selecting the Number of Neighbors
@@
Proposition D.12 (Minimization of Upper Bound).
@@
Combining Equations 75, 77 and 73 gives the result. ∎
See D.10
@@
The result now follows from Proposition D.12. ∎
D.5 Consistency of our Nearest Neighbor Method under Infill Asymptotics
In this section we prove Corollary 5.2, which establishes the consistency of SNN under infill asymptotics. The idea of the proof is to upper bound the $k$th order fill distance of the validation points in the test points to the fill distance of the validation points in $[0,1]^{d}$. This allows together with Theorem 5.1 and Proposition D.10 allows us to derive an upper bound on the error of our method in terms of the fill distance, from which the result follows.
D.5.1 Preliminary Lemma: Relating Fill Distances
Proposition D.13.
@@ D.5.1 Preliminary Lemma: Relating Fill Distances
D.5.2 Bound on Loss Depending on Fill Distance
@@
Corollary D.14 (Bound for Dense Validation Data and General Test Data).
@@
Choose $k=\min(\lceil(\gamma\tilde{\rho})^{-\frac{2d}{d+2}}\rceil,N^{\textup{{val}}})$ for some $\gamma\in(0,1/2)$ to be specified later. Because $\frac{2d}{d+2}<d$ and $\gamma\leq\frac{1}{2}$, this $k$ satisfies the conditions of Proposition D.13 so,
@@
We now apply Theorem 5.1 and Proposition D.10 together with this bound to conclude with probability $1-\delta$
@@
D.5.3 Consistency of Spatial Nearest Neighbors
We now restate and prove that the spatial nearest neighbor procedure we describe is consistent under infill asymptotics. This follows as a corollary of Corollary D.14 since the infill assumption means that for any fixed $\delta$, the upper bound in Corollary D.14 tends to zero with the fill distance.
@@ D.5.3 Consistency of Spatial Nearest Neighbors
Take $\delta=\min(1,r)$. Under infill asymptotics, this tends to zero because $r\leq C\tilde{\rho}$ and $\tilde{\rho}$ tends to 00 by assumption. Corollary D.14 (or more precisely Equation 96 which makes the depend of the bound on $\delta$ explicit), with probability at least $1-\delta$
@@
D.5.4 Computation of An Approximate Fill Distance
The fill distance can be computed exactly by computing the vertices of a $d$-dimensional Voronoi diagram with the validation points, then computing the maximum distance from each of these vertices to a point in the validation set. For problems in $1$ or $2$ dimensions, this is feasible even with a large number of validation points, because algorithms for computation of the Voronoi diagram can be done in nearly linear, $O(N^{\textup{{val}}}\log N^{\textup{{val}}})$ time in $1$ and $2$ dimensions [Okabe et al., 2000, Chapter 4]. In higher dimensions, the worst case complexity of algorithms for computing Voronoi diagrams can be $O((N^{\textup{{val}}})^{\lfloor d/2\rfloor}+N^{\textup{{val}}}\log N^{\textup{{val}}})$, and so for large number of points in more the computational cost can become quite high.
@@ D.5.4 Computation of An Approximate Fill Distance
D.6 Model Selection: Rates of Convergence of Spatial Nearest Neighbors
We now present an extended version of earlier discussion in Appendix C, as well as results on rates of convergence of spatial nearest neighbors that provides some support to claims regarding model selection. As a special case, we consider grid prediction (for example for assessing the global performance of a map constructed using a predictive method). We then discuss rates of convergence in the more general setting earlier addressed in Corollary D.14.
D.6.1 A Heuristic Discussion of Model Selection with Increasing Amounts of Data
@@
We will assume we are in the additive, homoskedastic error setting so that $Y=f(S,\chi(S))+\epsilon$. We will assume that the noise is bounded, and that squared loss is used. Finally, we will assume that training data is also generated following this process. We will also assume that $f$ is Lipschitz continuous. In this setting, minimax pointwise regression rates are $\theta((N^{\textup{{train}}})^{-\frac{1}{d+2}})$ (cf. Tsybakov [2008, Theorem 2.3, Corollary 2.2] in one-dimension under a fixed grid design and Tibshirani [2023, Examples 3.1, 3.2] for a multi-dimensional version with both fixed grid and random design. The latter assumes Gaussian noise instead of bounded noise). Ideally, we would like a method for performing model selection to be able to distinguish between a sequence of predictive methods converging at slower than the minimax optimal rate and a sequence of predictive methods converging at the minimax rate and to be able to reliably select between two sequences of predictive methods both of which are converging at the minimax rate, but with different constants.
@@
D.6.2 Rates of Convergence for Grid Prediction

### v3

@@ C.3.3 Inconsistency of Nearest Neighbors with Number of Neighbors Depending On Number of Validation Points
The idea is that either 1.) $(k_{n})_{n=1}^{\infty}$ has a bounded sub-sequence, in which case along this sub-sequence, the $\hat{R}_{N^{\textup{{val}}}}(h)$ can have a variance bounded below by 00, and so by proposition C.8 these estimators are bounded away from the test risk with fixed probability. Or 2.) the number of neighbors used tends to infinity, in which case we can find a sequence of data that accumulates more slowly around the test site, leading to many neighbors far from the point being used in the estimator, and therefore non-negligible bias.
@@
$\mathrm{Var}(\hat{R}_{N^{\textup{{val}}}})$ $=\mathrm{Var}\Big{(}\sum_{n=1}^{N^{\textup{{val}}}}w^{\textup{{NN}},k}|\epsilon_{n}|\Big{)}$ (47)
@@
Applying proposition C.8 to the random variable $|\hat{R}_{N^{\textup{{val}}}}(h)-R(h)|$ we conclude there exists an $\epsilon,\delta$ such that with probability at least $1-\delta$,
@@
C.3.4 Inconsistency of Blocked Spatial Validation
In what follows, we first review blocked spatial validation; then we prove it is inconsistent under infill asymptotics; and finally we provide illustrative examples that support its inconsistency (and the spatial consistency of our proposed method). The intuition is the same as in our previous holdout counterexamples: blocked spatial validation is inconsistent because the set of locations used for validation could be substantially different from the set of locations where we want to apply the model (the test locations).
Blocked Spatial Validation.
In blocked spatial validation, we have access to a collection of data, with spatial locations, that may be used for training and validation. All observations within a contiguous spatial area, typically a square (i.e., block), are reserved for validation. As an example, see the left plot in fig. 4, where we treat the horizontal and vertical dimensions as two spatial coordinates (e.g., latitude and longitude). We imagine that we originally have access to all of the training (pink circles) and validation (orange diamonds) data. We choose a particular spatial block (here, one sixteenth of the area of the interest) and reserve all of the data within it for validation. The estimate of test risk is the empirical average of the loss over the validation data. This estimate can be seen as a particular example of the holdout estimator of test risk. While we have already seen that the holdout need not be spatially consistent, our counterexample above (proposition 4.1) was not a spatially blocked example, so it still remains to show blocked spatial validation is inconsistent, as we do below.
Existing Work on Blocked Spatial Validation.
Blocked spatial validation is typically used as part of a cross-validation scheme, blocked spatial cross-validation (also sometimes referred to as spatial block cross-validation or simply block cross-validation) (Burman et al., 1994; Roberts et al., 2017). In blocked spatial cross-validation, the region of interest is partitioned into disjoint regions. In a particular fold of spatial cross-validation, data within a single region are reserved for validation data, and the data in the remaining regions serve as the training data for the fold. For instance, we can imagine partitioning the square of spatial locations into sixteen blocks. Figure 4 can then be interpreted as a single fold of blocked spatial cross-validation. As in sections 4 and A.1, we emphasize that, while cross-validation can be useful for data efficiency, even in i.i.d. settings it essentially estimates a different notion of risk than validation does (Bates et al., 2023). So we leave a careful study of blocked spatial cross-validation to future work and focus on evaluating blocked spatial validation presently.
Spatial (cross-)validation is commonly motivated by a desire to test the extrapolation quality of a prediction method (Roberts et al., 2017, Table 1). While the present paper does not examine the quality of validation methods for extrapolation tasks, we can still ask whether spatial validation passes the minimal check we propose in the present paper; that is, we ask whether spatial validation satisfies spatial consistency. In fact, we will shortly show it is spatially inconsistent.
Wadoux et al. (2021, Figure 1) and De Bruin et al. (2022, Figure 9) have previously used simulations to show empirically that spatial cross-validation can give biased estimates of map accuracy of a prediction method. Map accuracy is the loss between observed and predicted values of the response evaluated over a grid of prediction points in a particular region. This is essentially the same as the test risk we consider when the test point locations form a grid. A difference between our test risk with test point locations on a regular grid and map accuracy is that we treat the response as a random variable and average over this randomness as well, whereas map accuracy treats the test response values as fixed. However, no work has previously established the spatial consistency properties of spatial validation.
There is an intuitive understanding in some of the literature on estimating map accuracy that there may be a mismatch between the validation and test locations. Several recent works have tried to address the mismatch in the cross-validation setting, where validation locations in each fold can be chosen (Milà et al., 2022; Linnenbrink et al., 2023). (Conversely our only assumption in the present work is that validation locations arrive in a way that satisfies infill asymptotics.) These authors have suggested matching the distribution of pairwise distances between validation locations and the closest training locations to the distribution of pairwise distances between test locations and the closest training locations within each fold. We still expect these methods to exhibit bias in their estimates of test risk unless one is willing to make additional assumptions beyond the smoothness assumption in assumption 2.5.
We Show Blocked Spatial Validation is Inconsistent under Infill Asymptotics.
We will focus on cases where infill asymptotics is satisfied, and show that even in these particularly nice cases where validation data is available near the test point, the blocked spatial validation estimator need not give consistent estimates of the test risk. As in our other holdout counterexamples, the intuition is that the validation locations could be meaningfully different from the test locations, and getting more validation locations need not alleviate this mismatch.
See 4.2
We now construct such an example. In proposition C.10, we show that for this example (which has random validation locations), the fill distance tends to zero in probability and the holdout converges in probability to a different value from the test risk. The predictive method described is linear, and is therefore Lipschitz. And the response is a deterministic, Lipschitz function. Therefore, assumption 2.5 is also satisfied. Therefore, infill asymptotics is satisfied, and so proposition 4.2 follows from proposition C.10.
We prove proposition C.10 after describing the data-generating process, test distribution, and model that we will use.
Training Locations, Test Locations and Validation Locations. We generate the train locations on a $2$-dimensional regular grid on $[-0.5,0]\times[-0.5,5]$ with $5$ points in the first dimension and $11$ points in the second dimension, $S^{\text{train}}=\left\{-0.5,-0.4,-0.3,-0.2,-0.1\right\}\times\left\{-0.5,-0.4,-0.3,-0.2,-0.1,0,0.1,0.2,0.3,0.4,0.5\right\}$. We generate the validation locations uniformly at random on $[0,0.5]\times[-0.5,5]$. The test locations consist of a single point at $(0.15,0)$.
Covariates and Response. We consider a data-generating process without additional covariates. We assume the response is generated as $Y_{i}^{j}=|(S_{i}^{j})^{(1)}|_{1}$, for $j\in\{\text{train},\text{val},\text{test}\}$. In our analysis, we assume that there is no noise for simplicity.
Predictive Model. We consider ordinary least squares fit on the spatial location using the training data.
Blocked Spatial Validation Estimator. We use the holdout estimator in eq. 1. The training and validation data are generated to be contained in disjoint regions in space. The estimator is therefore an example of blocked spatial validation.
Proposition C.10.
[…]
Proof.
Because the training points all have first coordinate less than zero, $Y^{\text{train}}_{j}=(S^{\text{train}}_{j})^{(i)}$ for all $1\leq j\leq 55$. Therefore, the predictive model is $\hat{Y}(S)=-S^{(1)}$, as this has zero squared error on the training data, and is the unique linear function with this property. We can then compute the test loss as $|Y^{\textup{{test}}}_{1}-\hat{Y}((0.15,0))|=0.3$, proving the second part of the claim. By the strong law of large numbers, and because the holdout points are assumed to be independent and identically distributed, the holdout estimator converges almost surely to its expected value. We can calculate this expected value as,
[…]
Finally, the fill distance tends to 00 in probability by proposition 3.3, as the uniform distribution has non-zero density in a neighborhood containing the test point. ∎
Illustrative Simulation of Proposition C.10.
We simulate the process described above for number of validation points $N^{\textup{{val}}}\in\{250,500,1000,2000,4000,8000\}$. For each number of validation points, we resample the validation locations, the only random part of the process described above, 100 times; since the response at the validation points varies deterministically with the location in this simulation, the validation responses will update when the locations update. We compare our method, the blocking estimator described above, and the standard holdout (using all available validation data). We expect our method to be consistent, because the second half of the data is uniformly distributed on a region containing the test data, and so the fill distance tends to 00 with probability $1$ (proposition 3.3).
Figure 3: On the left: the training (pink circles), validation (orange diamonds), and test (blue triangle) locations for the blocked spatial validation counterexample described above. On the right: A comparison of our method and the holdout estimator using the spatial block on the dataset described above.
Figure 3 shows the train, validation and test locations. The plot on the right in fig. 3 supports the claim that the absolute error in risk estimation using the blocked holdout converges to $0.5-0.3=0.2$. And the error of our method is close to 00 as infill asymptotics holds.
Illustrative Simulation with More Complex Block Structure.
The example above is particularly simple with a single validation block and a single training block. We also consider a second simulation that more closely resembles the use of blocked spatial validation in practice while still satisfying an infill assumption. We generate all data on the unit square. For each seed, we select a block with side length $0.25\times 0.25$ that will only contain validation and test data. The remaining part of the unit square only contains test data so the data is spatially blocked into training and validation sets. The validation data are uniform on this sub-block. We sample the training data uniform on the remaining part of the unit square. We draw the test locations as a product of beta-distributions with parameter $0.5,0.5$ on the sub-square containing the validation data. As a result, the test data is clustered close to the edges of the square containing validation data. We use 1000 train points and 500 test points. We vary the number of the validation points in $\{250,500,1000,2000\}$. We generate the response as a Gaussian process on these locations with Matérn 5/2 kernel with lengthscale 0.1, and Gaussian noise with standard deviation $0.1$. We use truncated mean absolute error loss, truncated at $2$. We fit kernel ridge regression as the predictive method, using the kernel used to generate the data and the variance ($0.01$) as the regularization parameter. The training, validation and test locations are shown in fig. 4, left.
Figure 4: On the left: the training (pink circles), validation (orange diamonds) and test (blue triangle) locations for the blocked spatial validation counterexample described above. On the right: A comparison of our method and the holdout estimator using the spatial block on the dataset described above.
We resample the spatial locations, as well as the Gaussian process and noise at these spatial locations, as described above 100 times. In fig. 4, we compare the quality of estimation of blocked spatial validation and spatial nearest neighbors (ours) on these 100 seeds. We see that the holdout using blocking has substantially larger errors than our method (SNN). Also, while our method improves as the amount of validation data increases, as our theory suggests should happen since infill asymptotics is satisfied, the blocked spatial validation estimator using the holdout does not improve substantially even as we increase the amount of validation data available by an order of magnitude.
C.4 General Nearest Neighbor Bound and Selecting the Number of Neighbors
@@
Proposition C.11 (Minimization over Powers of Two).
@@
C.4.1 Preliminary Result: Hoeffding’s Inequality
@@
Lemma C.12 (Hoeffding’s Inequality, (Hoeffding, 1963, Theorem 2)).
@@
C.4.2 Proof of General Nearest Neighbor Risk Estimation Bound
We again recall and prove theorem 5.1. The idea is to apply triangle inequality to split the error into a bias term and a sampling error term. The sampling error term is then controlled with lemma C.12 since the loss is bounded. The bias term is controlled using assumption 2.5. See 5.1
@@ C.4.2 Proof of General Nearest Neighbor Risk Estimation Bound
The first term, $\tau_{1}$ is a bias term, while the second term, $\tau_{2}$ is a sum of $N^{\textup{{val}}}$ independent variables with expectation zero. Using assumption 2.1, we apply Hoeffding’s inequality (lemma C.12) to bound $\tau_{2}$: for any $\delta\in(0,1)$ with probability at least $1-\delta$
@@
Substituting this in eq. 63 and using eq. 62
@@
The result follows from combining eqs. 60, 61, 72 and 66. ∎
C.4.3 Proofs Related to Selecting the Number of Neighbors
@@
Proposition C.13 (Minimization of Upper Bound).
@@
Combining eqs. 77, 79 and 75 gives the result. ∎
See C.11
@@
The result now follows from proposition C.13. ∎
C.5 Consistency of our Nearest Neighbor Method under Infill Asymptotics
In this section we prove corollary 5.2, which establishes the consistency of SNN under infill asymptotics. The idea of the proof is to upper bound the $k$th order fill distance of the validation points in the test points to the fill distance of the validation points in $[0,1]^{d}$. This allows together with theorem 5.1 and proposition C.11 allows us to derive an upper bound on the error of our method in terms of the fill distance, from which the result follows.
C.5.1 Preliminary Lemma: Relating Fill Distances
Proposition C.14.
@@ C.5.1 Preliminary Lemma: Relating Fill Distances
C.5.2 Bound on Loss Depending on Fill Distance
@@
Corollary C.15 (Bound for Dense Validation Data and General Test Data).
@@
Choose $k=\min(\lceil(\gamma\tilde{\rho})^{-\frac{2d}{d+2}}\rceil,N^{\textup{{val}}})$ for some $\gamma\in(0,1/2)$ to be specified later. Because $\frac{2d}{d+2}<d$ and $\gamma\leq\frac{1}{2}$, this $k$ satisfies the conditions of proposition C.14 so,
@@
We now apply theorem 5.1 and proposition C.11 together with this bound to conclude with probability $1-\delta$
@@
C.5.3 Consistency of Spatial Nearest Neighbors
We now restate and prove that the spatial nearest neighbor procedure we describe is consistent under infill asymptotics. This follows as a corollary of corollary C.15 since the infill assumption means that for any fixed $\delta$, the upper bound in corollary C.15 tends to zero with the fill distance.
@@ C.5.3 Consistency of Spatial Nearest Neighbors
Take $\delta=\min(1,r)$. Under infill asymptotics, this tends to zero because $r\leq C\tilde{\rho}$ and $\tilde{\rho}$ tends to 00 by assumption. Corollary C.15 (or more precisely eq. 98 which makes the depend of the bound on $\delta$ explicit), with probability at least $1-\delta$
@@
C.5.4 Computation of An Approximate Fill Distance
The fill distance can be computed exactly by computing the vertices of a $d$-dimensional Voronoi diagram with the validation points, then computing the maximum distance from each of these vertices to a point in the validation set. For problems in $1$ or $2$ dimensions, this is feasible even with a large number of validation points, because algorithms for computation of the Voronoi diagram can be done in nearly linear, $O(N^{\textup{{val}}}\log N^{\textup{{val}}})$ time in $1$ and $2$ dimensions (Okabe et al., 2000, Chapter 4). In higher dimensions, the worst case complexity of algorithms for computing Voronoi diagrams can be $O((N^{\textup{{val}}})^{\lfloor d/2\rfloor}+N^{\textup{{val}}}\log N^{\textup{{val}}})$, and so for large number of points in more the computational cost can become quite high.
@@ C.5.4 Computation of An Approximate Fill Distance
C.6 Model Selection: Rates of Convergence of Spatial Nearest Neighbors
We now present an extended version of earlier discussion in appendix B, as well as results on rates of convergence of spatial nearest neighbors that provides some support to claims regarding model selection. As a special case, we consider grid prediction (for example for assessing the global performance of a map constructed using a predictive method). We then discuss rates of convergence in the more general setting earlier addressed in corollary C.15.
C.6.1 A Heuristic Discussion of Model Selection with Increasing Amounts of Data
@@
We will assume we are in the additive, homoskedastic error setting so that $Y=f(S,\chi(S))+\epsilon$. We will assume that the noise is bounded, and that squared loss is used. Finally, we will assume that training data is also generated following this process. We will also assume that $f$ is Lipschitz continuous. In this setting, minimax pointwise regression rates are $\theta((N^{\textup{{tr}}})^{-\frac{1}{d+2}})$ (cf. Tsybakov (2008, Theorem 2.3, Corollary 2.2) in one-dimension under a fixed grid design and Tibshirani (2023, Examples 3.1, 3.2) for a multi-dimensional version with both fixed grid and random design. The latter assumes Gaussian noise instead of bounded noise). Ideally, we would like a method for performing model selection to be able to distinguish between a sequence of predictive methods converging at slower than the minimax optimal rate and a sequence of predictive methods converging at the minimax rate and to be able to reliably select between two sequences of predictive methods both of which are converging at the minimax rate, but with different constants.
@@
C.6.2 Rates of Convergence for Grid Prediction